# Pay Attention to Real World Perturbations! Natural Robustness Evaluation in Machine Reading Comprehension

### v1

@@ 6 Dealing With Natural Perturbations
In this section, we provide an initial exploration of methods to defend against natural perturbations, focusing on encoder-only models and SQuAD datasets. Expanding to other datasets and architectures could be explored in future work. To enhance model robustness, we conduct adversarial training by identifying six encoder-only model architectures that already exhibit the highest robustness to natural perturbations in their respective categories (except albert-xxlarge-v2 on Nat_V2_Challenge), and presenting them with both original training data and the generated naturally perturbed training examples. We extract the entire Wikipedia revision histories for the articles in the original SQuAD training set, and then obtain (with questions) and (with questions) perturbed contexts to augment the original SQuAD 1.1 and SQuAD 2.0 training set, respectively, using the methodology described in Section 3. Table 4 compares the performance of these models on Nat_V1_Challenge and Nat_V2_Challenge, before and after retraining.
@@ 7 Conclusion
We acknowledge several limitations in our work: (1) Our perturbation framework constructs natural perturbations from Wikipedia edit history and therefore only works with Wikipedia-based benchmarks. Since the phenomenon of natural perturbations is by no means limited to Wikipedia and can occur in any kind of text that evolves over time, future work should explore alternative methods to generate natural perturbations for non-Wikipedia MRC datasets. (2) As training data augmentation has a relatively limited impact, further research is needed to develop better techniques for improving the robustness of encoder-only models to natural perturbations and to investigate the relationship between robustness to natural and synthetic perturbations. Enhancing the robustness of LLMs is also an important direction for future work. (3) Potential data contamination may affect our findings on LLM evaluation. Investigating its extent and impact on robustness evaluation will be a focus of our future research efforts.

### v2

@@ 6 Dealing With Natural Perturbations
In this section, we provide an initial exploration of methods to defend against natural perturbations. To enhance encoder-only model robustness, we first conduct adversarial training by identifying six encoder-only model architectures that already exhibit the highest robustness to natural perturbations in their respective categories (except albert-xxlarge-v2 on Nat_V2_Challenge), and presenting them with both original training data and the generated naturally perturbed training examples. We extract the entire Wikipedia revision histories for the articles in the original SQuAD training set, and then obtain (with questions) and (with questions) perturbed contexts to augment the original SQuAD 1.1 and SQuAD 2.0 training set, respectively, using the methodology described in Section 3. Table 4 compares the performance of these models on Nat_V1_Challenge and Nat_V2_Challenge, before and after retraining.
@@
Similarly, for the LLMs, we adopt a few-shot prompting approach by including both the original MRC instance and its naturally or synthetically perturbed counterpart as demonstrations, and assess how model performance and robustness change compared to the zero-shot setting (see Table 5 and Table 6). A total of two original–perturbed instance pairs are used, with the original samples taken from the SQuAD training set. Although not widely observed, in certain cases, in-context demonstrations can improve an LLM’s resilience to natural perturbations, regardless of whether natural or synthetic perturbed examples are demonstrated. This phenomenon is particularly evident in models such as Llama-3.2-3B-Instruct, OLMo-7B-0724-Instruct and Falcon3-10B-Instruct. However, it can also have detrimental effects, further decreasing LLM robustness and resulting in a performance decline on both the clean and naturally perturbed test sets.
| LLM | Nat_V1_Challenge | Nat_V2_Challenge | ||||
| Orig.Pert. | IM Drop | zero-shot | Orig.Pert. | IM Drop | zero-shot | |
| Gemma 2-2b-IT | / | / | ||||
| Gemma 2-9b-IT | / | / | ||||
| Llama-3.1-8B-Instruct | / | / | ||||
| Llama-3.2-3B-Instruct | / | / | ||||
| Mistral-7B-Instruct-v0.2 | / | / | ||||
| OLMo-7B-0724-Instruct | / | / | ||||
| Qwen2.5-3B-Instruct | / | / | ||||
| Qwen2.5-7B-Instruct | / | / | ||||
| Qwen2.5-14B-Instruct | / | / | ||||
| Falcon3-7B-Instruct | / | / | ||||
| Falcon3-10B-Instruct | / | / | ||||
| deepseek-llm-7b-chat | / | / | ||||
| LLM | Nat_V1_Challenge | Nat_V2_Challenge | ||||
| Orig.Pert. | IM Drop | zero-shot | Orig.Pert. | IM Drop | zero-shot | |
| Gemma 2-2b-IT | / | / | ||||
| Gemma 2-9b-IT | / | / | ||||
| Llama-3.1-8B-Instruct | / | / | ||||
| Llama-3.2-3B-Instruct | / | / | ||||
| Mistral-7B-Instruct-v0.2 | / | / | ||||
| OLMo-7B-0724-Instruct | / | / | ||||
| Qwen2.5-3B-Instruct | / | / | ||||
| Qwen2.5-7B-Instruct | / | / | ||||
| Qwen2.5-14B-Instruct | / | / | ||||
| Falcon3-7B-Instruct | / | / | ||||
| Falcon3-10B-Instruct | / | / | ||||
| deepseek-llm-7b-chat | / | / | ||||
@@ 7 Conclusion
We acknowledge several limitations in our work: (1) Our perturbation framework constructs natural perturbations from Wikipedia edit history and therefore only works with Wikipedia-based benchmarks. Since the phenomenon of natural perturbations is by no means limited to Wikipedia and can occur in any kind of text that evolves over time, future work should explore alternative methods to generate natural perturbations for non-Wikipedia MRC datasets. (2) As training data augmentation and in-context demonstration have a relatively limited impact, further research is needed to develop better techniques for improving the robustness of both encoder-only models and LLMs to natural perturbations, and to investigate the relationship between robustness to natural and synthetic perturbations. (3) Potential benchmark contamination may affect our findings on LLM evaluation. Investigating its extent and impact on LLM performance and robustness evaluation will be a focus of our future research efforts.